# On the inclusion of non-concurrent controls in platform trials with a futility interim analysis

### v1

@@ Abstract
The analysis of platform trials can be enhanced by utilizing non-concurrent controls. Since including this data might also introduce bias in the treatment effect estimators if time trends are present, methods for incorporating non-concurrent controls adjusting for time have been proposed. However, so far their behavior has not been systematically investigated in platform trials that include interim analyses. To evaluate the impact of a futility interim analysis in trials utilizing non-concurrent controls, we consider a platform trial featuring two experimental arms and a shared control, with the second experimental arm entering later. We focus on a frequentist regression model that uses non-concurrent controls to estimate the treatment effect of the second arm and adjusts for time using a step function to account for temporal changes. We show that performing a futility interim analysis in Arm 1 may introduce bias in the point estimation of the effect in Arm 2, if the regression model is used without adjustment, and investigate how the marginal bias and bias conditional on the first arm continuing after the interim depend on different trial design parameters. Moreover, we propose a new estimator of the treatment effect in Arm 2, aiming to eliminate the bias introduced by both the interim analysis in Arm 1 and the time trends, and evaluate its performance in a simulation study. The newly proposed estimator is shown to substantially reduce the bias and type I error rate inflation while leading to power gains compared to an analysis using only concurrent controls.
@@ 1 Introduction
Several approaches aiming to utilize non-concurrent controls while ensuring valid statistical inference have recently been proposed. In particular, Lee and Wason [8] and Bofill Roig et al. [9] considered incorporating the NCC data using a regression model with categorical time adjustment, including the periods in the trial as a fixed effect, where periods are defined as time intervals bounded by any experimental arm entering or leaving the platform. This approach was investigated in a simple platform trial with two treatment arms and it was shown that this model yields unbiased treatment effect estimates if the time trends are equal across all arms and additive on the model scale [9]. Krotka et al. [10] examined the performance of this model in more complex platform trials with an arbitrary number of treatment arms and extended the frequentist methodology for incorporating non-concurrent controls by proposing more flexible methods, such as spline regression or mixed models. Another frequentist method proposed by Marschner and Schou [11] uses a network meta-analysis approach to analyze the platform trial as a network of direct randomized comparisons and indirect non-randomized comparisons. Methods based on propensity score weighting for incorporating NCC in platform trials with time trends under the framework of causal inference were considered by Guo et al. [12]. Among Bayesian approaches considered for utilizing NCC data are the Bayesian Time Machine [13] and the meta-analytic-predictive (MAP) prior approach [14, 15]. The Time Machine approach uses a Bayesian hierarchical model to smooth the control response over calendar time intervals. The MAP prior approach borrows data from the non-concurrent periods to obtain the prior distribution for the control response in the concurrent periods, while accounting for the between-trial heterogeneity. Several frequentist and Bayesian methods for incorporating NCC data were recently compared in a comprehensive simulation study [16]. Interim analyses in platform trials were discussed by Greenstreet et al. [17] in the context of a multi-stage design that allows for the late entry of new arms in a pre-planned way. They proposed an approach for performing interim analyses pre-specified in the planning phase, while controlling the family-wise error rate. However, this approach does not incorporate non-concurrent controls, nor does it account for potential time trends. Including interim analyses in platform trial designs that aim to utilize non-concurrent controls has not yet been discussed in the literature.
@@
Even though the bias in treatment effect estimates in classical group sequential trials has been well studied, the impact of conducting an interim analysis on the effect estimation in a platform trial utilizing non-concurrent controls in the analysis has not yet been investigated. In this work, we aim to examine how, in platform trials that evaluate multiple experimental treatments, the effect estimation in late-entering arms is affected by interim results in earlier arms. For simplicity, we focus on a two-arm platform trial, where the second arm enters the trial later on, and an interim analysis for Arm 1 is performed at the time when Arm 2 is added. Such a timing of the interim analysis enables monitoring of the ongoing trial at the time of addition of the new arm [17]. We examine the previously proposed model-based approach for including non-concurrent controls that includes time as a fixed effect [8, 9]. Focusing on platform trials with continuous endpoints, we describe how the weight of non-concurrent controls included in the estimation of the treatment effect of Arm 2 depends on the interim result in Arm 1, and show that applying the current regression model, as considered by Lee and Wason [8] and Bofill Roig et al. [9], in group sequential platform trials leads to biased treatment effect estimators and type I error rate inflation. Following the idea from Whitehead [22], we propose a mean adjusted estimator to mitigate this bias. We derive the analytical expression for the bias introduced due to the interim analysis enabling futility stopping, and investigate the performance of the newly proposed estimator in terms of the bias, root mean squared error, type I error rate, and power in a simulation study. The extensions to efficacy stopping are discussed in the Discussion Section.
@@ 2 Design Setting
Consider a platform trial with $2$ treatment arms (indexed by $k=1,2$) and a common control group ($k=0$), where Arm 1 starts at the beginning of the trial and Arm 2 enters the ongoing trial later on. Both arms finish the trial at the same time. The trial is split into two periods, divided by the time point where Arm 2 joins the platform. For Arm 1, an interim analysis is performed, with the possibility of dropping this arm due to futility. It is assumed that this interim analysis is conducted at the time point of adding Arm 2. The second arm is only assessed in the final analysis at the end of the trial, after the total sample size is reached. In this final analysis, we aim to compare the efficacy of Arm 2 with the shared control group, i.e., including the NCC data for this arm. In this work, we consider platform trials with a continuous endpoint. The considered design is illustrated in Figure 1.
@@ 2.1 Regression model with period adjustment in trials without interim analysis
$E(y_{j})=\eta_{0}+\sum_{k=1,2}\theta_{k}\cdot I(k_{j}=k)+\tau\cdot I(s_{j}=2)$ (1)
@@
$\tilde{y}_{02}=(1-\varrho)\cdot\bar{y}_{02}+\varrho\cdot[\bar{y}_{01}+\bar{y}_{12}-\bar{y}_{11}]$ (3)
@@
3 Bias introduced by futility interim analysis in regression-based estimators
@@
Suppose that an interim analysis of Arm 1 is performed at the time when Arm 2 is added to the trial. This interim analysis is based on a one-sided z-test comparing the mean responses between Arm 1 and control in period 1, i.e., testing $H_{0}:\theta_{1}=0$ vs $H_{1}:\theta_{1}>0$. We consider stopping for futility if the one-sided $p$-value from the z-test is larger than a futility bound $\alpha_{1}$. Note that when including an interim analysis for Arm 1, the treatment effect estimator for Arm 2 depends on the interim decision. If Arm 1 continues in the trial, the estimator given by (2) depends on the non-concurrent control data, since by (4) $\varrho>0$. However, if Arm 1 stops at the interim analysis, $n_{12}=0$ and therefore the weight $\varrho$ of the non-concurrent control data becomes zero and the model does not utilize the NCC data. In this case, the treatment effect estimate of Arm 2 is $\tilde{\theta}_{2}=\bar{y}_{22}-\bar{y}_{02}$. Hence, the extent of borrowing from the NCC data and thus the model-based treatment effect estimator in the final analysis for Arm 2 is stochastically dependent on the interim treatment effect estimate for Arm 1. As a result, it is also dependent on the non-concurrent control data. Below, we derive the marginal and conditional bias of the model-based treatment effect estimator for Arm 2 for the platform trial design in Figure 1.
@@
Assume there is a stepwise time trend with strength $\lambda$, i.e., the mean response increases by $\lambda$ at the beginning of period 2. Hence, the sample means in period 1 $\bar{y}_{01}$ and $\bar{y}_{11}$ follow normal distributions with means $\mu_{0}$ and $\mu_{1}$, respectively. In period 2, the sample means $\bar{y}_{02}$, $\bar{y}_{12}$, and $\bar{y}_{22}$ follow normal distributions with means $\mu_{0}+\lambda$, $\mu_{1}+\lambda$, and $\mu_{2}+\lambda$, respectively. Note that the treatment effects of arms 1 and 2, denoted by $\theta_{1}=\mu_{1}-\mu_{0}$ and $\theta_{2}=\mu_{2}-\mu_{0}$, respectively, are constant across the whole trial, since the time trend affects all arms equally. Denote by $Z_{11}=(\bar{y}_{11}-\bar{y}_{01})/(\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}})$ the interim Z-statistic for Arm 1. The arm is stopped for futility if the interim p-value $p_{11}=1-\Phi(Z_{11})$ is larger than a futility boundary $\alpha_{1}$, hence if $Z_{11}<\Phi^{-1}(1-\alpha_{1})=c_{1}$. The probability to stop Arm 1 at the interim analysis is $\text{I\kern-1.49994ptP}(Z_{11}<c_{1})=\Phi(\gamma)$, with $\gamma=c_{1}-\theta_{1}/(\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}})$. Note that under the null hypothesis for Arm 1, this probability simplifies to $1-\alpha_{1}$.
@@
As shown in Appendix A, $\tilde{\theta}_{2}$ is biased. Its marginal bias –that is, the expected bias when averaging over interim outcomes (Arm 1 stopping in the interim analysis and continuing to period 2)– is given by:
@@
$E[\tilde{\theta}_{2}-\theta_{2}]=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\phi\left(\gamma\right)$ (5)
@@
where $\phi(x)$ is the probability density function of the standard normal distribution evaluated at $x$. Note that the bias does not depend on the strength of the time trend $\lambda$. It is not caused by the time trend but arises from the interim analysis. Moreover, conditional on the event that Arm 1 stops, $\tilde{\theta}_{2}$ is unbiased. This follows since, in this case, the treatment effect estimate in Arm 2 does not depend on the data from period 1. Hence, the bias of $\tilde{\theta}_{2}$ conditional on the event that Arm 1 continues is given by the marginal bias (5) divided by the probability that Arm 1 continues:
$E[\tilde{\theta}_{2}-\theta_{2}|Z_{11}\geq c_{1}]=\frac{E[\tilde{\theta}_{2}-\theta_{2}]}{1-\Phi\left(\gamma\right)}$ (6)
@@
The extent of this bias depends on: the weight of the NCC data when Arm 1 continues, $\varrho$ (i.e., the preplanned sample sizes in Arm 1 and control); the futility boundary, $\alpha_{1}$; and the standardized effect size in Arm 1, $\theta_{1}/(\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}})$. The model-based estimator is biased under both the null and alternative hypotheses for Arm 1. However, the bias does not depend on the treatment effect in Arm 2, $\theta_{2}$. Note that under the null hypothesis for Arm 1 ($\theta_{1}=0$), (6) simplifies to:
@@
$E[\tilde{\theta}_{2}-\theta_{2}|Z_{11}\geq c_{1}]=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(c_{1}\right)}{\alpha_{1}}$ (7)
@@
Figure 2: Marginal and conditional bias of the unadjusted model-based treatment effect estimator for Arm 2 (2) when varying different design parameters. In all cases, no treatment effect for Arm 1 is assumed ($\theta_{1}=0$), and a unit variance $\sigma^{2}=1$ in each arm is used.
A) Varying the futility bound $\alpha_{1}$. Sample sizes are set to 150 per arm and period.
@@
B) Varying the sample sizes ratio between period 1 and period 2 $r=n_{01}/n_{02}=n_{11}/n_{12}$. Sample sizes per arm in period 2 are fixed ($n_{02}=n_{12}=150$). Futility bound $\alpha_{1}=0.5$ is used.
@@
C) Varying the allocation ratio in Arm 1 vs control $a=n_{11}/n_{01}=n_{12}/n_{02}$. The sample size in the control arm is fixed to 150 in each period ($n_{01}=n_{02}=150$). Futility bound $\alpha_{1}=0.5$ is used.
@@
Figure 2 shows the marginal (5) and conditional (6) bias when varying different design parameters of the considered platform trial. Figure 2-A, shows the bias as a function of the futility bound $\alpha_{1}$, when $\theta_{1}=0$. While the conditional bias is decreasing in $\alpha_{1}$, the marginal bias is maximized for $\alpha_{1}=0.5$, i.e., when Arm 1 has a 50% probability to continue after the interim analysis. Note that for $\alpha_{1}=0$ and $\alpha_{1}=1$, the treatment effect estimator $\tilde{\theta}_{2}$ is marginally unbiased, since these futility bounds result either in always using the separate analysis (since Arm 1 always stops) or always using the model-based approach (since Arm 1 always continues), respectively. The behavior of the bias with respect to the sample size ratio between period 1 and period 2 is shown in Figure 2-B. In particular, the sample sizes per arm in period 2 are set to 150, and the ratio $r=n_{01}/n_{02}=n_{11}/n_{12}$ is varied. Both, the marginal and conditional bias are maximized for $r=1$, i.e. when the periods are equally sized. This is because for smaller period 1 sample sizes ($r<1$), the weight of the period 1 data from Arm 1 in the Arm 2 treatment effect estimate is smaller. On the other hand, for larger period 1 sample sizes ($r>1$), the period 1 estimates have smaller standard errors and therefore their conditional bias (conditional on the event that the trial continues) is smaller. The impact of the allocation ratio between Arm 1 and control on the bias is illustrated in Figure 2-C. Specifically, the sample size in the control arm is fixed in each period ($n_{01}=n_{02}=150$), and the sample size in Arm 1 is varied using the allocation ratio $a=n_{11}/n_{01}=n_{12}/n_{02}$. Note that the sample sizes per period are equal in each arm. Both marginal and conditional bias increase with the sample size in Arm 1 relative to the control sample size.

### v2

# On the inclusion of non-concurrent controls in platform trials with an interim analysis

@@ Abstract
The analysis of platform trials can be enhanced by utilizing non-concurrent controls. Since including this data might also introduce bias in the treatment effect estimators if time trends are present, methods for incorporating non-concurrent controls adjusting for time have been proposed. However, so far their behavior has not been systematically investigated in platform trials that include interim analyses. To evaluate the impact of an interim analysis in trials utilizing non-concurrent controls, we consider a platform trial featuring two experimental arms and a shared control, with the second experimental arm entering later. We focus on a frequentist regression model that uses non-concurrent controls to estimate the treatment effect of the second arm and adjusts for time using a step function to account for temporal changes. We show that performing an interim analysis in Arm 1 may introduce bias in the point estimation of the effect in Arm 2, if the regression model is used without adjustment, and investigate how the marginal bias and bias conditional on the first arm continuing after the interim depend on different trial design parameters. Moreover, we propose a new estimator of the treatment effect in Arm 2, aiming to eliminate the bias introduced by both the interim analysis in Arm 1 and the time trends, and evaluate its performance in a simulation study. The newly proposed estimator is shown to substantially reduce the bias and type I error rate inflation while leading to power gains compared to an analysis using only concurrent controls.
@@ 1 Introduction
Several approaches aiming to utilize non-concurrent controls while ensuring valid statistical inference have recently been proposed. In particular, Lee and Wason [18] and Bofill Roig et al. [3] considered incorporating the NCC data using a regression model with categorical time adjustment, including the periods in the trial as a fixed effect, where periods are defined as time intervals bounded by any experimental arm entering or leaving the platform. This approach was investigated in a simple platform trial with two treatment arms and it was shown that this model yields unbiased treatment effect estimates if the time trends are equal across all arms and additive on the model scale [3]. Krotka et al. [17] examined the performance of this model in more complex platform trials with an arbitrary number of treatment arms and extended the frequentist methodology for incorporating non-concurrent controls by proposing more flexible methods, such as spline regression or mixed models. Another frequentist method proposed by Marschner and Schou [20] uses a network meta-analysis approach to analyze the platform trial as a network of direct randomized comparisons and indirect non-randomized comparisons. Methods based on propensity score weighting for incorporating NCC in platform trials with time trends under the framework of causal inference were considered by Guo et al. [15]. Among Bayesian approaches considered for utilizing NCC data are the Bayesian Time Machine [26] and the meta-analytic-predictive (MAP) prior approach [27, 31]. The Time Machine approach uses a Bayesian hierarchical model to smooth the control response over calendar time intervals. The MAP prior approach borrows data from the non-concurrent periods to obtain the prior distribution for the control response in the concurrent periods, while accounting for the between-trial heterogeneity. Several frequentist and Bayesian methods for incorporating NCC data were recently compared in a comprehensive simulation study [24].
Platform trials involve multiple hypothesis tests because several experimental treatments are investigated within a single trial. In addition, multiple endpoints or patient subgroups may be evaluated [6, 23]. In this setting, one may consider control of the platformwise type I error rate, defined as the probability of falsely rejecting at least one null hypothesis across the entire platform trial. This corresponds to the familywise error rate for the full set of hypotheses tested in the platform. However, it has been argued that no multiplicity adjustment is required for inferentially independent tests, such as for treatments with different mechanisms of action [16]. In that case, the pairwise error rate, that is, the probability of falsely rejecting the null hypothesis for a particular arm, needs to be controlled. Nevertheless, control of the familywise error rate at the substudy level remains necessary, when multiple doses, endpoints or subgroups are tested for a particular treatment.
Implementing a multiple testing procedure that controls the type I error rate at the platform level is challenging because, at the start of a platform trial, neither the total number of hypotheses nor the hypotheses themselves is typically pre-specified. Furthermore, results from the different arms become available only sequentially over time. Therefore, online multiple testing procedures have been proposed to control either the false discovery rate, defined as the expected proportion of false rejections among all rejections, or the family-wise error rate, defined as the probability of at least one false rejection within a family of hypotheses [34, 23, 12, 13, 10]. In such settings, false discovery rate control is of particular interest because it is less conservative than family-wise error rate control and may therefore provide greater power when many hypotheses are tested over time. Interim analyses represent an additional source of multiplicity in group-sequential platform trials. Zehetmayer et al. [34] proposed several methods for false discovery rate control in this setting and investigated their performance in a simulation study. Interim analyses in platform trials were also discussed by Greenstreet et al. [12] in the context of a multi-stage design that allows late entry of new arms in a pre-planned manner. Assuming known entry times of the experimental arms and focusing mainly on settings without time trends, they proposed an approach for conducting interim analyses specified at the design stage while controlling the family-wise error rate. Related work by the same authors further considered pre-planned multi-stage platform trial designs in which multiple superior treatments may be identified [13]. However, the approaches of Greenstreet et al. [12, 13] do not incorporate non-concurrent controls. Including interim analyses in platform trial designs that aim to use non-concurrent controls has not yet been discussed in the literature.
@@
Even though the bias in treatment effect estimates in classical group sequential trials has been well studied, the impact of conducting an interim analysis on the effect estimation in a platform trial utilizing non-concurrent controls in the analysis has not yet been investigated. In this work, we aim to examine how, in platform trials that evaluate multiple experimental treatments, the effect estimation in late-entering arms is affected by interim results in earlier arms. For simplicity, we focus on a two-arm platform trial, where the second arm enters the trial later on, and an interim analysis for Arm 1 is performed at the time when Arm 2 is added. Such a timing of the interim analysis enables monitoring of the ongoing trial at the time of addition of the new arm [12]. We examine the previously proposed model-based approach for including non-concurrent controls that includes time as a fixed effect [18, 3]. Focusing on platform trials with continuous endpoints, we describe how the weight of non-concurrent controls included in the estimation of the treatment effect of Arm 2 depends on the interim result in Arm 1, and show that applying the current regression model, as considered by Lee and Wason [18] and Bofill Roig et al. [3], in group sequential platform trials leads to biased treatment effect estimators and type I error rate inflation. Following the idea from Whitehead [32], we propose a mean adjusted estimator to mitigate this bias. We derive the analytical expression for the bias introduced due to the interim analysis enabling futility and efficacy stopping, and investigate the performance of the newly proposed estimator in terms of the bias, root mean squared error, type I error rate, and power in a simulation study.
@@ 2 Design Setting
Consider a platform trial with $2$ treatment arms (indexed by $k=1,2$) and a common control group ($k=0$), where Arm 1 starts at the beginning of the trial, and Arm 2 enters the ongoing trial later on. Both arms finish the trial at the same time. The trial is split into two periods, divided by the time point where Arm 2 joins the platform. For Arm 1, an interim analysis is performed, with the possibility of stopping this arm due to futility or efficacy. It is assumed that this interim analysis is conducted at the time point of adding Arm 2. The second arm is only assessed in the final analysis at the end of the trial, after the total sample size is reached. In this final analysis, we aim to compare the efficacy of Arm 2 with the shared control group, i.e., including the NCC data for this arm. Direct pooling of NCC and CC data is statistically invalid in the presence of time trends; when unaccounted for, time trends can introduce severe bias into the treatment effect estimator [7, 5, 29]. In this work, we therefore utilize a frequentist model-based approach to incorporate the NCC data into the analysis of Arm 2 while adjusting for time trends [3, 18]. The considered design is illustrated in Figure 1.
@@ 2.1 Regression model with period adjustment in trials without interim analysis
$E(y_{j})=\eta_{0}+\sum_{k=1,2}\theta_{k}\cdot I(k_{j}=k)+\tau\cdot I(s_{j}=2)$
@@
$\tilde{y}_{02}=(1-\varrho)\cdot\bar{y}_{02}+\varrho\cdot[\bar{y}_{01}+\bar{y}_{12}-\bar{y}_{11}]$
@@
3 Bias introduced by an interim analysis in regression-based estimators
@@
Suppose that an interim analysis of Arm 1 is performed at the time when Arm 2 is added to the trial. This interim analysis is based on a z-test comparing the mean responses between Arm 1 and control in period 1, i.e., testing $H_{0}:\theta_{1}=0$ vs $H_{1}:\theta_{1}>0$. We consider stopping for futility if the $p$-value from the z-test is larger than a futility bound $\alpha_{F}$ ($0\leq\alpha_{F}\leq 1$), and stopping for efficacy if it is below an efficacy boundary $\alpha_{E}$ ($0\leq\alpha_{E}\leq 1$, $\alpha_{E}<\alpha_{F}$). Note that when including an interim analysis for Arm 1, the treatment effect estimator for Arm 2 depends on the interim decision. If Arm 1 continues in the trial, the estimator given by (1) depends on the non-concurrent control data, since by (2) $\varrho>0$. However, if Arm 1 stops at the interim analysis, $n_{12}=0$ and therefore the weight $\varrho$ of the non-concurrent control data becomes zero and the model does not utilize the NCC data. In this case, the treatment effect estimate of Arm 2 is based only on the CC data: $\tilde{\theta}_{2}=\bar{y}_{22}-\bar{y}_{02}$. Thus, whether non-concurrent control data are borrowed in the final estimator for Arm 2 is determined by the interim decision for Arm 1. Since this interim decision depends on the interim treatment effect estimate for Arm 1, it also depends on the non-concurrent control data. This dependence induces bias. Below, we derive the marginal and conditional bias of the model-based treatment effect estimator for Arm 2 for the platform trial design shown in Figure 1.
@@
Assume patient $j$ enters the trial at time $t_{j}$, and is enrolled in period $s_{j}$. Further, assume a time trend defined by a function $f(t_{j})$ of the patient entry time $t_{j}$ ($j=1,\ldots,N$), which is the same for all arms and additive on the model scale, as described in Bofill Roig et al. [3]. For each period $s$, define $m_{s}:=E[f(t_{j})\mid s_{j}=s]$ and assume that $E[f(t_{j})\mid k_{j}=k,s_{j}=s]=m_{s}$ for all arms $k$ that are open for randomization in period $s$. We additionally assume that the variances of the outcomes $y_{j}$ are equal across periods. This approximately holds when the strength of the time trend is small relative to the response variance conditional on time. Hence, the sample means in period 1 $\bar{y}_{01}$ and $\bar{y}_{11}$ follow normal distributions with means $\mu_{0}+m_{1}$ and $\mu_{1}+m_{1}$, respectively. In period 2, the sample means $\bar{y}_{02}$, $\bar{y}_{12}$, and $\bar{y}_{22}$ follow normal distributions with means $\mu_{0}+m_{2}$, $\mu_{1}+m_{2}$, and $\mu_{2}+m_{2}$, respectively. Note that the treatment effects of arms 1 and 2, denoted by $\theta_{1}=\mu_{1}-\mu_{0}$ and $\theta_{2}=\mu_{2}-\mu_{0}$, respectively, are constant across the whole trial, since the time trend affects all arms equally. Denote by $Z_{11}=(\bar{y}_{11}-\bar{y}_{01})/(\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}})$ the interim Z-statistic for Arm 1. The arm is stopped for futility if the interim p-value $p_{11}=1-\Phi(Z_{11})$ is larger than a futility boundary $\alpha_{F}$, hence if $Z_{11}<\Phi^{-1}(1-\alpha_{F})=c_{F}$. Similarly, the arm stops for efficacy if the p-value is below an efficacy boundary $\alpha_{E}$, i.e., $Z_{11}>\Phi^{-1}(1-\alpha_{E})=c_{E}$, where $c_{E}>c_{F}$. Under the true treatment effect $\theta_{1}$, the probability to stop Arm 1 at the interim analysis is $\text{I\kern-1.49994ptP}(Z_{11}<c_{F})+\text{I\kern-1.49994ptP}(Z_{11}>c_{E})=\Phi(c_{F}-\delta)+\Phi(\delta-c_{E})$, with $\delta=\theta_{1}/(\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}})$. Note that under the null hypothesis for Arm 1, this probability simplifies to $(1-\alpha_{F})+\alpha_{E}$.
@@
Since the inclusion of NCC data now stochastically depends on the interim outcome, $\tilde{\theta}_{2}$ is biased. As shown in Appendix A, its marginal bias –that is, the expected bias when averaging over interim outcomes (Arm 1 stopping in the interim analysis and continuing to period 2)– is given by:
@@
$E[\tilde{\theta}_{2}-\theta_{2}]=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\left(\phi\left(c_{F}-\delta\right)-\phi\left(c_{E}-\delta\right)\right)$ (3)
@@
where $\phi(x)$ is the probability density function of the standard normal distribution evaluated at $x$. Note that the bias does not depend on the time trend function $f(t_{j})$, as long as the randomization is independent of the entry times $t_{j}$ in each period, since the bias is not caused by the time trend but arises from the interim analysis. Moreover, conditional on the event that Arm 1 stops, $\tilde{\theta}_{2}$ is unbiased. This follows since, in this case, the treatment effect estimate in Arm 2 does not depend on the data from period 1. Hence, the bias of $\tilde{\theta}_{2}$ conditional on the event that Arm 1 continues is given by the marginal bias (3) divided by the probability that Arm 1 continues:
$E[\tilde{\theta}_{2}-\theta_{2}|c_{F}\leq Z_{11}\leq c_{E}]=\frac{E[\tilde{\theta}_{2}-\theta_{2}]}{\Phi(c_{E}-\delta)-\Phi(c_{F}-\delta)}$ (4)
@@
where $\Phi(x)$ is the cumulative distribution function of the standard normal distribution evaluated at $x$. The extent of this bias depends on: the weight of the NCC data when Arm 1 continues, $\varrho$ (i.e., the preplanned sample sizes in Arm 1 and control); the futility and efficacy boundaries, $\alpha_{F}$ and $\alpha_{E}$; and the standardized effect size in Arm 1, $\delta=\theta_{1}/(\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}})$. The model-based estimator is biased under both the null and alternative hypotheses for Arm 1. However, the bias does not depend on the treatment effect in Arm 2, $\theta_{2}$. Note that under the null hypothesis for Arm 1 ($\theta_{1}=0$), (4) simplifies to:
@@
$E[\tilde{\theta}_{2}-\theta_{2}|c_{F}\leq Z_{11}\leq c_{E}]=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(c_{F}\right)-\phi\left(c_{E}\right)}{\alpha_{F}-\alpha_{E}}$ (5)
@@
Figure 2: Marginal and conditional bias of the unadjusted model-based treatment effect estimator for Arm 2 (1) when varying different design parameters. In all cases, no treatment effect for Arm 1 is assumed ($\theta_{1}=0$), and a unit variance $\sigma^{2}=1$ in each arm is used. In the first row, only stopping for futility is applied ($\alpha_{E}=0$ in all cases), while in the second row, only efficacy stopping is used ($\alpha_{F}=1$ in all cases).
A) Varying the futility bound $\alpha_{F}$ and $\alpha_{E}$. Sample sizes are set to 150 per arm and period.
@@
B) Varying the sample sizes ratio between period 1 and period 2 $r=n_{01}/n_{02}=n_{11}/n_{12}$. Sample sizes per arm in period 2 are fixed ($n_{02}=n_{12}=150$). Stopping bounds $\alpha_{F}=0.5$ (first row) and $\alpha_{E}=0.5$ (second row) are used.
@@
C) Varying the allocation ratio in Arm 1 vs control $a=n_{11}/n_{01}=n_{12}/n_{02}$. The sample size in the control arm is fixed to 150 in each period ($n_{01}=n_{02}=150$). Stopping bounds $\alpha_{F}=0.5$ (first row) and $\alpha_{E}=0.5$ (second row) are used.
@@
Figure 2 shows the marginal (3) and conditional (4) bias when varying different design parameters of the considered platform trial. If only futility stopping is applied (first row in Figure 2), the unadjusted model-based estimator is positively biased. This happens due to the underestimation of the mean response in the non-concurrent controls in cases where Arm 1 does not stop at the interim analysis, which results in the overestimation of the treatment effect in Arm 2. In contrast, with interim analyses allowing early stopping for efficacy only (second row in Figure 2), the unadjusted model-based estimator is, by symmetry, negatively biased. This occurs because the mean response in the non-concurrent controls is overestimated when Arm 1 continues, leading to underestimation of the treatment effect in Arm 2. Figure 2-A shows the bias as a function of the futility bound $\alpha_{F}$ (first row), or of the efficacy bound $\alpha_{E}$ (second row), when $\theta_{1}=0$. The absolute conditional bias increases with decreasing probability for Arm 1 to continue, i.e., with decreasing $\alpha_{F}$, and with increasing $\alpha_{E}$. The absolute marginal bias is maximized for $\alpha_{F}=0.5$ or $\alpha_{E}=0.5$, i.e., when Arm 1 has a 50% probability to continue after the interim analysis. Note that for $\alpha_{F}=0$ and $\alpha_{F}=1$, as well as $\alpha_{E}=1$ and $\alpha_{E}=0$, the treatment effect estimator $\tilde{\theta}_{2}$ is marginally unbiased, since these bounds result either in always using the separate analysis (since Arm 1 always stops) or always using the model-based approach (since Arm 1 always continues). The behavior of the bias with respect to the sample size ratio between period 1 and period 2 is shown in Figure 2-B. In particular, the sample sizes per arm in period 2 are set to 150, and the ratio $r=n_{01}/n_{02}=n_{11}/n_{12}$ is varied. Both the absolute marginal and conditional bias are maximized for $r=1$, i.e., when the periods are equally sized. This is because for smaller period 1 sample sizes ($r<1$), the weight of the period 1 data from Arm 1 in the Arm 2 treatment effect estimate is smaller. On the other hand, for larger period 1 sample sizes ($r>1$), the period 1 estimates have smaller standard errors and therefore their conditional bias (conditional on the event that the trial continues) is smaller. The impact of the allocation ratio between Arm 1 and control on the bias is illustrated in Figure 2-C. Specifically, the sample size in the control arm is fixed in each period ($n_{01}=n_{02}=150$), and the sample size in Arm 1 is varied using the allocation ratio $a=n_{11}/n_{01}=n_{12}/n_{02}$. Note that the sample sizes per period are equal in each arm. Both marginal and conditional bias increase in their absolute value with the sample size in Arm 1 relative to the control sample size.
When both futility and efficacy stopping rules are applied, the marginal bias of the model-based estimator is influenced by both stopping rules, and its sign depends on the treatment effect in Arm 1 (see Figure 3, grey line). For very small treatment effects in Arm 1, early stopping occurs predominantly for futility, resulting in a positive marginal bias in $\tilde{\theta}_{2}$. For very large treatment effects in Arm 1, early stopping occurs predominantly for efficacy, resulting in a negative marginal bias in $\tilde{\theta}_{2}$. For intermediate effect sizes, the opposing impacts of futility and efficacy stopping may partially offset one another, reducing the magnitude of the bias.
Figure 3: Conditional (left figure) and marginal (right figure) bias of the unadjusted model-based treatment effect estimator for Arm 2 (1) when varying the treatment effect in Arm 1 ($\theta_{1}$). In all cases, sample sizes of 150 in each arm and period are assumed, and unit variance $\sigma^{2}=1$ in each arm is used. Different lines show different stopping rules: futility stopping only ($\alpha_{F}=0.5$, $\alpha_{E}=0$, red line), efficacy stopping only ($\alpha_{F}=1$, $\alpha_{E}=0.00258$, blue line), and stopping for both futility and efficacy ($\alpha_{F}=0.5$, $\alpha_{E}=0.00264$, grey line). The chosen efficacy boundary $\alpha_{E}$ corresponds to the O’Brien-Fleming boundary using binding futility stopping at $\alpha_{F}$, assuming a significance level of 0.025 in the final analysis. Note that for better readability, the y-axis scale is different in each figure.